X = o,1,2,3 and P(X)= 2k,3k,13k,2k.

how is this problem done?

a random variable X has a probability distribution as follows done?

What is K and why is its value .05   

Total probability = 1
2k+3k+13k+2k= 1
k= .05
p(x<2)= 2k + 3k = 5k = 5*.05=.25
Answer is B 0.25
Answer #1

here as we know that probability sum of all events is 1

therefore here sum of all probability events =P(X=0)+P(X=1)+P(X=2)+PX=3)=1

2k+3k+13k+2k=1

20k=1

k=1/20=0.05

for P(x<2)=P(X=0)+P(X=1)=2k+3k =5k=5*(1/20)=0.25

(please revert if required any clarification)